# The Role of Entanglement in Quantum Reservoir Computing with Coupled Kerr Nonlinear Oscillators

### v2

@@ Abstract
Quantum Reservoir Computing (QRC) uses quantum dynamics to efficiently process temporal data. In this work, we investigate a QRC framework based on two coupled Kerr nonlinear oscillators, a system well-suited for time-series prediction tasks due to its complex nonlinear interactions and potentially high-dimensional state space. We explore how its performance in forecasting both linear and nonlinear time-series depends on key physical parameters: input drive strength, Kerr nonlinearity, and oscillator coupling, and analyze the role of entanglement in improving the reservoir’s computational performance, focusing on its effect on predicting non-trivial time series. Using logarithmic negativity to quantify entanglement and normalized root mean square error (NRMSE) to evaluate predictive accuracy, our results suggest that entanglement provides a computational advantage on average—up to a threshold in the input frequency—that persists under some levels of dissipation and dephasing. In particular, we find that higher dissipation rates can enhance performance. While the entanglement advantage manifests as improvements in both average and worst-case performance, it does not lead to improvements in the best-case error. These findings contribute to the broader understanding of quantum reservoirs for high performance, efficient quantum machine learning and time-series forecasting.
@@ I Introduction
To address these open questions, we investigate a QRC framework based on two coupled Kerr nonlinear oscillators. We employ Hamiltonian encoding, consistent with previous oscillator-based QRC implementations [15, 9, 19], rather than the state injection techniques often utilized in spin-network memory capacity studies [22]. This strategy is particularly advantageous for forecasting as it circumvents the experimental overhead of repeated state preparation and enables highly nonlinear feature mapping. Furthermore, it avoids the control overheads associated with resetting the reservoir’s memory [27], ensuring that entanglement is not totally destroyed each time. Moreover, we do time-series forecasting both on linear multi-frequency signals and nonlinear NARMA time-series, which is substantially different than memory tasks because forecasting necessitates learning the signal’s underlying dynamics rather than merely storing past inputs. Crucially, this approach uncovers a behavior distinct from spin-network QRC [22]: instead of a monotonic performance gain at high frequencies, we observe a frequency threshold, below which entanglement provides the computational advantage.
@@ II.2 Learning and Simulation
$NRMSE=\frac{1}{y_{max}-y_{min}}\sqrt{\frac{\sum_{i}^{N}(y_{i}-\tilde{y}_{i})^{2}}{N}}$ (7)
@@ III Results
Figure 2: Time evolution of entanglement (top row) and corresponding time-series prediction performance (bottom row) at different input frequencies $f$. Top: Behavior of logarithmic negativity in time during testing and training. To ensure the reservoir reflects stable dynamics rather than initial transients, we use the portion of the evolution after the vertical dashed line for training and testing. Bottom: Examples of input sequences at four different frequency scales and The corresponding predictions of the reservoir at coupling rate of $g=0.9$ , input strengths of $\epsilon_{a}=\epsilon_{b}=3$, nonlinearities of $K_{a}=K_{b}=1$ and dissipation rates of $\kappa_{a}=\kappa_{b}=0.1$ in a rotating frame. The performance gets worse as the input frequency increases.
@@ III.1 Input Strength
By focusing on the entanglement (right panel of Fig. 3), we observe that it increases up to $\epsilon\simeq 2$ for all frequencies. Beyond this point, entanglement decreases, leading to a peak in entanglement that corresponds to the first NRMSE dip, especially for $f=0.8$ and $f=1.5$. It should be noted that while the best performance is not achieved with maximum entanglement, it does occur at a moderate degree of entanglement. Therefore, a certain amount of entanglement appears to be advantageous.
@@
Figure 3: NRMSE (left) and entanglement (right) versus input strength $\epsilon$ for different input frequencies, with fixed parameters $K=1$, $\kappa_{a}=\kappa_{b}=0.1$, $g=0.9$, and $\kappa_{\phi}=0$. Left: While the first dips are observed at $\epsilon\simeq 2$, NRMSE decreases with increasing input strength up to $\epsilon\simeq 5$, after which it saturates or worsens depending on the input frequency. Right: Logarithmic negativity increases with $\epsilon$ until a peak near $\epsilon\simeq 2$, beyond which it declines with a lower slope. Best-case errors correspond to moderate entanglement for frequencies below the threshold (e.g., $f=2.5$).
@@ III.2 Kerr Nonlinearity
The relation is somewhat more subtle here. The NRMSE for $f=0.8$ and $f=1.5$ (green and orange points in Fig.4) fluctuates up to $K\simeq 2$. Beyond this point, the NRMSE starts increasing. For $f=0.05$, it increases monotonically, and for $f=2.5$, no clear trend is observed (left panel of Fig.4).
By focusing on the entanglement (right panel of Fig. 4), we observe that it fluctuates up to $K\simeq 1$, which makes it hard to infer any advantage up to this point. Beyond this point, entanglement decreases while the error increases, which could be interpreted as a hint of entanglement advantage for all frequencies below the threshold, i.e., $f=2.5$. Overall, we can say that best-case errors correspond to moderate entanglement for frequencies below the threshold (e.g., $f=2.5$).
@@
Figure 4: NRMSE (left) and Entanglement (right) versus Kerr nonlinearity $K$ for different input frequencies, with fixed parameters $\epsilon=3$, $\kappa_{a}=\kappa_{b}=0.1$, $g=0.9$, and $\kappa_{\phi}=0$. Left: NRMSE fluctuates up to $K\simeq 2$ for $f=0.8$ and $f=1.5$, and increases thereafter. For $f=0.05$, it increases monotonically, while no clear trend is observed for $f=2.5$. Right: Entanglement fluctuates up to $K\simeq 1$, beyond which it decreases. Best-case errors correspond to moderate entanglement for frequencies below the threshold (e.g., $f=2.5$).
@@
Figure 5: NRMSE (left) and Entanglement (right) versus coupling constant $g$ for different input frequencies, with fixed parameters $\epsilon=3$, $\kappa_{a}=\kappa_{b}=0.1$, $K=1$, and $\kappa_{\phi}=0$. Left: NRMSE for $f=0.8$ and $f=1.5$ fluctuates up to $g\simeq 1.4$, beyond which it increases. For $f=0.05$, NRMSE increases monotonically, and no clear trend is observed for $f=2.5$. Right: Entanglement peaks near $g\simeq 0.9$. Beyond this point, entanglement decreases. Best-case errors (third points from the left on the left panel) correspond to moderate entanglement (the same points for the right panel) for frequencies below the threshold (e.g., $f=2.5$).
@@ III.3 Coupling Constant
By focusing on the entanglement (right panel of Fig. 5), we observe that it peaks at $g\simeq 0.9$, which suggests no clear correlation with NRMSE. Beyond this point, entanglement decreases while the error increases, which could be interpreted as a hint of entanglement advantage for all frequencies below the threshold. Best-case errors correspond to moderate entanglement for frequencies below the threshold (e.g., $f=2.5$).
III.4 Entanglement Advantage
We now consolidate the results from the sections III.1, III.2, and III.3 by performing a binned analysis: the data are grouped into multiple bins with equal sizes based on entanglement levels, and a summary statistic of the NRMSE is computed for each bin. This allows us to observe the overall relationship between entanglement and predictive performance.
@@ III.4 Entanglement Advantage
Our analysis reveals a non-monotonic relationship between entanglement and predictive performance. As shown in Fig. 6, we first consider the minimum NRMSE in each bin (left panel), which shows no discernible pattern as entanglement varies. Next, we observe the average predictive performance: moderate levels of entanglement are associated with improved average predictive performance, with NRMSE reaching a minimum within a specific frequency regime of the input. However, for frequency scale f=0.05f=0.05(purple curve in Fig. 6), error saturates for higher entanglement, whereas a dip is observed for other cases. Beyond this regime, increasing the input frequency leads to a breakdown in reservoir memory and a decline in average performance, despite sustained or even increasing levels of entanglement, as the task becomes too difficult for the predictor in this setup. A potential reason for this could be the low excitation number limit that was implemented to reduce the artifacts caused by the cutoff dimension (II.4 and appendix B). Finally, the maximum NRMSE in each bin (right panel) exhibits a similar but slightly stronger relation to entanglement compared to the mean NRMSE.
Refer to caption Refer to caption Refer to caption
Figure 6: Prediction error (NRMSE) vs. entanglement in the case of zero dephasing ($\kappa_{\phi}=0$) and a dissipation rate of $\kappa_{a}=\kappa_{b}=0.1$, for different input frequencies. The data from figs. 3, 4 and 5 are combined and binned based on entanglement levels, quantified via logarithmic negativity, with vertical error bars representing the standard deviation within each bin. Left: Minimum NRMSE within each entanglement bin. No consistent correlation is observed. Middle: Mean NRMSE across all configurations in each bin. This captures the overall performance trend with respect to entanglement. Moderate entanglement improves performance on average. There is an optimum level of entanglement. However, a threshold in the frequency is observed at $f=2.5$ where no trend exists. Right: Maximum NRMSE per entanglement bin. This reflects the worst-performing cases within each bin. A stronger relation is seen here: higher entanglement reduces the worst-case error more than the average error, but again, a frequency threshold is observed. Note: Vertical error bars in all panels represent the standard deviation of the entire data distribution within that entanglement bin. They are included in the Min/Max panels to illustrate the dispersion of the performance data associated with that specific level of entanglement, rather than the uncertainty of the extrema values themselves.
@@ III.5 Role of Dissipation
To examine the impact of dissipation, we repeat the data generation process and binned analysis described earlier in the presence of different dissipation rates, $\kappa_{a}=\kappa_{b}=\kappa$ and compute the average entanglement and NRMSE to produce the same type of plot as the middle panel of Fig. 6, now for different $\kappa$ (see Fig. 9 in Appendix A). It is worth mentioning that our system does not reach a steady state in the absence of dissipation and therefore does not exhibit steady-state entanglement; hence, we do not consider the zero-dissipation case. To better analyze the effect of dissipation, we extract the optimal points (those with minimum error) and plot their corresponding NRMSE as a function of dissipation for different input frequencies, resulting in Fig. 7.
As shown in Fig. 7, a higher dissipation level is better for frequencies below the frequency threshold up to a point where it saturates. By looking at Fig. 9 in Appendix A one can see that realizing an entanglement advantage requires sufficient dissipation, with slightly higher values offering a consistent advantage across all frequency scales below the frequency threshold.
Figure 7: Minimum NRMSE at optimal points vs. input frequency for different dissipation rates where $\kappa_{\phi}=0$. These optimal points correspond to the minimum average NRMSE in Fig. 6 and Fig. 9 in Appendix A. Higher dissipation improves performance for all frequencies below the frequency threshold, up to a point in dissipation where it tends to saturate.
@@ III.6 Role of Dephasing
As one can see in Fig. 8 and Fig. 10 in Appendix A, zero dephasing is the optimal regime both for overall performance and the entanglement advantage (because at high dephasing we do not have much entanglement). However, focusing on Fig. 8 reveals that for very low frequency (e.g. $f=0.05$) some amount of dephasing could be beneficial.
Figure 8: Minimum NRMSE at optimal points vs. input frequency for different dephasing rates where $\kappa_{a}=\kappa_{b}=0.1$. These optimal points correspond to the minimum average NRMSE in Fig. 6 and Fig. 10 in Appendix A. Zero or very small dephasing yields the best performance in most frequency regimes. However, at very low frequency ($f=0.05$), a small amount of dephasing slightly improves prediction performance.
@@ III.7 Nonlinear Prediction Task
To verify the generality of our findings beyond the linear forecasting task, we extended our analysis to a more complex nonlinear prediction task. We applied the 20th-order NARMA kernel to the multi-frequency input signal to generate a highly nonlinear target trajectory. As detailed in Appendix C, our analysis confirms that the key trends observed in the linear task remain robust. Specifically, we find that moderate entanglement continues to be linked to the optimal predictive performance. Furthermore, the frequency threshold effect persists, where the entanglement advantage is lost at higher input frequencies. Interestingly, while dissipation remains beneficial up to a saturation point , we observe that dephasing is strictly detrimental to performance in this nonlinear regime, highlighting the increased sensitivity of nonlinear tasks to phase coherence. Full details and parameter sweeps for this task are provided in Appendix C.
@@ IV Discussion and Conclusion
Our findings indicate that within the various system configurations analyzed, entanglement is indeed linked to improvement in both average and worst-case predictive performance, albeit only within a certain range of operational parameters and up to a threshold in the input frequency. It was noted that optimal average performance, measured in terms of NRMSE, generally takes place when the reservoir maintains a moderate level of entanglement rather than the highest, with the worst-case performance following a similar trend. This effect was most evident for lower input frequency scales, where the temporal structure of the input could be effectively retained and processed by the reservoir dynamics. This kind of relation between quantumness and learning capabilities is consistent with other recent findings in the literature [29, 14, 22], which suggest that excessive “quantumness” may saturate the reservoir or introduce dynamical instability, ultimately hampering learning. However, unlike the spin-network results in [22] which observed an increasing advantage at high frequencies, our coupled-oscillator system exhibits a distinct frequency threshold, which might be related to the excitation number cut-off and the corresponding limit on the average excitation number that we imposed.
Notably, our results demonstrate that the observed connection between entanglement and performance persists under realistic conditions with dissipation and dephasing (please see Appendix A). We showed that increasing dissipation can actually improve performance, up to a point that it saturates. Likewise, we found that the absence of dephasing is generally optimal for performance and entanglement (please see Appendix A); a small amount of dephasing can be tolerated (and at very low input frequency, it even provided a slight benefit).
While our results demonstrate a clear link between entanglement and prediction accuracy, we note that establishing strict causality in physical reservoirs is challenging. However, the robustness of this relation—persisting across different parameters, noise levels, and tasks (including the nonlinear NARMA time-series prediction task, as shown in Appendix C)—suggests that entanglement is linked to the reservoir’s forecasting power in the quantum regime.
@@
Nevertheless, as described above, we were able to find a connection between entanglement and performance. This finding could potentially strengthen the broader hypothesis that quantum resources can be beneficial in machine learning frameworks, provided that system parameters and noise levels are appropriately balanced. Looking forward, extending this analysis to larger oscillator networks would allow for exploration of more complex spatial and temporal entanglement structures. It would also be valuable to consider alternative quantum resource measures and output processing techniques, including non-linear readout layers or adaptive hybrid quantum-classical schemes.

### v3

@@ Abstract
Quantum Reservoir Computing (QRC) uses quantum dynamics to efficiently process temporal data. In this work, we investigate a QRC framework based on two coupled Kerr nonlinear oscillators, a system well-suited for time-series prediction tasks due to its complex nonlinear interactions and potentially high-dimensional state space. We explore how its performance in forecasting both linear and nonlinear time-series depends on key physical parameters: input drive strength, Kerr nonlinearity, and oscillator coupling, and analyze the role of entanglement in improving the reservoir’s computational performance, focusing on its effect on predicting non-trivial time series. Using logarithmic negativity to quantify entanglement and normalized root mean square error (NRMSE) to evaluate predictive accuracy, individual parameter sweeps show that optimal performance occurs at moderate but non-zero entanglement. Furthermore, an aggregated binned analysis reveals that this moderate entanglement is consistently associated with the optimal average predictive performance across the parameter space, an observation that persists up to a threshold in the input frequency. This relationship persists under some levels of dissipation and dephasing. In particular, we find that higher dissipation rates can enhance performance. These findings contribute to the broader understanding of quantum reservoirs for high performance, efficient quantum machine learning and time-series forecasting.
@@ I Introduction
To address these open questions, we investigate a QRC framework based on two coupled Kerr nonlinear oscillators. We employ Hamiltonian encoding, consistent with previous oscillator-based QRC implementations [15, 9, 19], rather than the state injection techniques often utilized in spin-network memory capacity studies [22]. This strategy is particularly advantageous for forecasting as it circumvents the experimental overhead of repeated state preparation and enables highly nonlinear feature mapping. Furthermore, it avoids the control overheads associated with resetting the reservoir’s memory [27], ensuring that entanglement is not totally destroyed each time. Moreover, we do time-series forecasting both on linear multi-frequency signals and nonlinear NARMA time-series, which is substantially different than memory tasks because forecasting necessitates learning the signal’s underlying dynamics rather than merely storing past inputs. Our approach employs both single-parameter sweeps and a binned analysis on the corresponding aggregated data. Through these methods, we uncover that moderate levels of entanglement coincide with optimal forecasting performance, up to a certain threshold in the input frequency. No such threshold was observed in the case of spin-network QRC [22].
@@ II.2 Learning and Simulation
$NRMSE=\frac{1}{y_{max}-y_{min}}\sqrt{\frac{\sum_{i=1}^{N}(y_{i}-\tilde{y}_{i})^{2}}{N}}$ (7)
@@ III Results
Figure 2: Time evolution of entanglement (top row) and corresponding time-series prediction performance (bottom row) at different input frequencies $f$. Top: Behavior of logarithmic negativity in time during testing and training. To ensure the reservoir reflects stable dynamics rather than initial transients, we use the portion of the evolution after the vertical dashed line for training and testing. Bottom: Examples of input sequences at four different frequency scales and the corresponding predictions of the reservoir at coupling rate of $g=0.9$, input strengths of $\epsilon_{a}=\epsilon_{b}=3$, nonlinearities of $K_{a}=K_{b}=1$, and dissipation rates of $\kappa_{a}=\kappa_{b}=0.1$ in a rotating frame. The performance gets worse as the input frequency increases.
@@ III.1 Input Strength
By focusing on the entanglement (right panel of Fig. 3), we observe that it increases up to $\epsilon\simeq 2$ for all frequencies. Beyond this point, entanglement decreases, leading to a peak in entanglement. Crucially, near-zero input strength yields the absolute highest prediction errors across all frequencies. Tracing the absolute minimum NRMSE for each frequency (marked by red triangles) reveals that the optimal predictive basin coincides with robust, moderate but non-zero entanglement levels (e.g., ranging between $0.11$ and $0.17$ for $f\leq 1.5$). Therefore, while the best performance is not achieved with the maximum entanglement, it coincides with a moderate level of entanglement.
@@
Figure 3: NRMSE (left) and entanglement (right) versus input strength $\epsilon$ for different input frequencies, with fixed parameters $K=1$, $\kappa_{a}=\kappa_{b}=0.1$, $g=0.9$, and $\kappa_{\phi}=0$. Red triangles denote the specific data points that yield the absolute minimum NRMSE for each frequency. Left: Near-zero input strength yields the worst prediction error. NRMSE decreases up to an optimal basin before saturating or worsening. Right: Logarithmic negativity increases with $\epsilon$ until a peak near $\epsilon\simeq 2$, beyond which it declines. By mapping the minimum NRMSE points (red triangles) to the entanglement panel, one can see that optimal forecasting performance coincides with moderate but non-zero levels of entanglement.
@@ III.2 Kerr Nonlinearity
The NRMSE for $f=0.8$ and $f=1.5$ (green and orange points in Fig. 4) fluctuates, reaching minimum values near $K\simeq 2$ before increasing. For $f=0.05$, the minimum error occurs at very low nonlinearity, after which it increases monotonically, and for $f=2.5$, no clear trend is observed. By explicitly mapping the global minimum NRMSE for each frequency onto the entanglement plot (indicated by red triangles), we observe a consistent pattern. The best-case prediction errors for $f=0.05$, $f=0.8$, and $f=1.5$ do not occur in the zero-entanglement regime. Instead, these optimal points coincide tightly with a moderate entanglement range spanning from approximately $0.12$ to $0.14$. Beyond this optimal region ($K\simeq 2$), entanglement generally decreases while the prediction error increases.
@@
Figure 4: NRMSE (left) and Entanglement (right) versus Kerr nonlinearity $K$ for different input frequencies, with fixed parameters $\epsilon=3$, $\kappa_{a}=\kappa_{b}=0.1$, $g=0.9$, and $\kappa_{\phi}=0$. Red triangles mark the data points corresponding to the absolute minimum NRMSE for each frequency below the threshold. Left: NRMSE fluctuates up to $K\simeq 2$ for $f=0.8$ and $f=1.5$ and increases thereafter. Right: Entanglement fluctuates up to $K\simeq 1$, beyond which it decreases. Tracing the optimal performance points (red triangles) confirms that the minimum prediction errors coincide with a moderate but non-zero level of entanglement.
@@
Figure 5: NRMSE (left) and Entanglement (right) versus coupling constant $g$ for different input frequencies, with fixed parameters $\epsilon=3$, $\kappa_{a}=\kappa_{b}=0.1$, $K=1$, and $\kappa_{\phi}=0$. Red triangles highlight the specific parameter instances yielding the lowest NRMSE for each valid frequency. Left: NRMSE fluctuates before eventually increasing at higher coupling strengths. Right: Entanglement peaks near $g\simeq 0.9$, beyond which it decreases. Notably, while the optimal coupling strength is small, the corresponding entanglement is non-zero. One can see in this figure that, once again, optimal forecasting performance does not occur in the zero-entanglement state, aligning instead with a moderate entanglement regime.
@@ III.3 Coupling Constant
By focusing on the entanglement (right panel of Fig. 5), we observe that it peaks at $g\simeq 0.9$. By examining the regime of minimum prediction error (indicated by the red triangles in Fig. 5), We observe that while intermediate frequencies achieve minimum error at weak coupling ($g\simeq 0.1$), the right panel demonstrates that the entanglement at these points is not zero, but has risen to moderate but non-zero levels ($E\simeq 0.06$ to $0.11$). Across all valid frequencies, the best-case error coincides with the moderate but non-zero entanglement regime.
In summary, across the preceding parameter sweeps, we consistently observe that the lowest prediction error (NRMSE) occurs at moderate but strictly non-zero entanglement, suggesting a positive role for entanglement in these predictive tasks.
III.4 Overall Entanglement-Performance Relationship
To complement the preceding single-parameter sweeps, which focused on identifying the regimes that yield optimal performance, we now analyze the average predictive performance across the combined data. We perform a binned analysis by grouping the data from sections III.1, III.2, and III.3 into equal-sized bins based on their entanglement levels and computing the mean NRMSE for each. This complementary approach allows us to observe the broader, overall relationship between entanglement and the reservoir’s forecasting capability.
@@ III.4 Overall Entanglement-Performance Relationship
Our analysis reveals a non-monotonic relationship between entanglement and predictive performance. As shown in Fig. 6, we observe that moderate but non-zero levels of entanglement are associated with improved average predictive performance, with NRMSE reaching a minimum, within a specific frequency regime of the input. However, for frequency scale $f=0.05$ (purple curve in Fig. 6), error saturates for higher entanglement, whereas a dip is observed for other frequencies below $f=2.5$. Beyond this regime, increasing the input frequency leads to a breakdown in reservoir memory and a decline in average performance, despite sustained or even increasing levels of entanglement, as the task becomes too difficult for the predictor in this setup. A potential reason for this could be the low excitation number limit that was implemented to reduce the artifacts caused by the cutoff dimension (II.4 and appendix B).
Figure 6: Prediction error (NRMSE) versus entanglement across different input frequencies, with
$\kappa_{\phi}=0$ and $\kappa_{a}=\kappa_{b}=0.1$. The data from figs. 3, 4 and 5 are combined and binned according to entanglement levels (quantified via logarithmic negativity). The plotted mean NRMSE for each bin captures the overall relationship between entanglement and performance, where vertical error bars denote the intra-bin standard deviation. The results demonstrate that a moderate but non-zero level of entanglement improves prediction accuracy. However, this trend vanishes above a frequency threshold($f=2.5$), beyond which entanglement yields no observable advantage.
@@ III.5 Role of Dissipation
To examine the impact of dissipation, we repeat the data generation process and binned analysis described earlier in the presence of different dissipation rates, $\kappa_{a}=\kappa_{b}=\kappa$ and compute the average entanglement and NRMSE to produce the same type of plot as in Fig. 6, now for different $\kappa$ (see Fig. 9 in Appendix A). It is worth mentioning that our system does not reach a steady state in the absence of dissipation and therefore does not exhibit steady-state entanglement; hence, we do not consider the zero-dissipation case. To better analyze the effect of dissipation, we extract the optimal points (those with minimum error) and plot their corresponding NRMSE as a function of dissipation for different input frequencies, resulting in Fig. 7.
As shown in Fig. 7, a higher dissipation level is better for frequencies below the frequency threshold up to a point where it saturates. By looking at Fig. 9, Appendix A, one can see that observing an entanglement-performance relationship requires sufficient dissipation, with slightly higher values offering a consistent advantage across all frequency scales below the frequency threshold.
Figure 7: Optimum mean NRMSE versus dissipation rate for different input frequencies, with $\kappa_{\phi}=0$. These data points correspond to the optimum values extracted from the mean NRMSE traces in Fig. 6 and Fig. 9 in Appendix A (error bars show the same standard deviations as Fig.
6 and Fig. 9). Higher dissipation improves performance for all frequencies below the frequency threshold, up to a point in dissipation where it tends to saturate.
@@ III.6 Role of Dephasing
As one can see in Fig. 8 and Fig. 10 in Appendix A, zero dephasing is the optimal regime both for overall performance and maintaining the entanglement-performance relationship (because at high dephasing we do not have much entanglement). However, focusing on Fig. 8 reveals that for very low frequency (e.g. $f=0.05$), some amount of dephasing could be beneficial.
Figure 8: Optimum mean NRMSE versus dephasing rate for different input frequencies, with $\kappa_{a}=\kappa_{b}=0.1$. these data points correspond to the optimum values extracted from the mean NRMSE traces in Fig.
6 and Fig. 10 in Appendix A (error bars show the same standard deviations as Fig. 6 and Fig. 10). Zero or very small dephasing yields the best performance in most frequency regimes. However, at very low frequency ($f=0.05$), a small amount of dephasing seems to slightly improve prediction performance.
@@ III.7 Nonlinear Prediction Task
To verify the generality of our findings beyond the linear forecasting task, we extended our analysis to a more complex nonlinear prediction task. We applied the 20th-order NARMA kernel to the multi-frequency input signal to generate a highly nonlinear target trajectory. As detailed in Appendix C, our analysis confirms that the key trends observed in the linear task remain robust. Specifically, we find that moderate entanglement continues to be linked to the optimal predictive performance. Furthermore, the frequency threshold effect persists, where the relation between entanglement and performance is lost at higher input frequencies. While dissipation remains beneficial up to a saturation point, we observe that dephasing is strictly detrimental to performance in this nonlinear regime, highlighting the increased sensitivity of nonlinear tasks to phase coherence. Full details and parameter sweeps for this task are provided in Appendix C.
@@ IV Discussion and Conclusion
Our findings indicate that within the various system configurations analyzed, moderate but non-zero entanglement coincides with the best predictive performance, albeit only within a certain range of dephasing rate and up to a threshold in the input frequency. It was noted that optimal average performance, measured in terms of NRMSE, generally takes place when the reservoir maintains a moderate level of entanglement rather than the highest, with the worst-case performance following a similar trend. This effect was most evident for lower input frequency scales, where the temporal structure of the input could be effectively retained and processed by the reservoir dynamics. This kind of relation between quantumness and learning capabilities is consistent with other recent findings in the literature [29, 14, 22], which suggest that excessive “quantumness” may saturate the reservoir or introduce dynamical instability, ultimately hampering learning. However, unlike the spin-network results in [22] which observed an increasing advantage at high frequencies, our coupled-oscillator system exhibits a distinct frequency threshold, which might be related to the excitation number cut-off and the corresponding limit on the average excitation number that we imposed.
Notably, our results demonstrate that the observed relationship between moderate entanglement and performance persists under realistic conditions with dissipation and dephasing (please see Appendix A). We showed that increasing dissipation can actually improve performance, up to a point that it saturates. Likewise, we found that the absence of dephasing is generally optimal for performance and entanglement (please see Appendix A); a small amount of dephasing can be tolerated (and at very low input frequency, it even provided a slight benefit).
@@
While our results exhibit a consistent relationship between entanglement and prediction accuracy, we note that establishing strict causality in physical reservoirs is challenging. Because our methodology relies on one-dimensional parameter sweeps and binned statistical analysis, we must be cautious in claiming a definitive causal “advantage”. However, the robustness of this relation—persisting across different parameters, noise levels, and tasks (including the nonlinear NARMA time-series prediction task, as shown in Appendix C)—suggests that moderate entanglement may play a positive role in the reservoir’s forecasting power in the quantum regime. Moreover, this finding lends some support to the broader hypothesis that quantum resources can be beneficial in machine learning frameworks, provided that system parameters and noise levels are appropriately balanced. Looking forward, extending this analysis to larger oscillator networks would allow for exploration of more complex spatial and temporal entanglement structures. It would also be valuable to consider alternative quantum resource measures and output processing techniques, including non-linear readout layers or adaptive hybrid quantum-classical schemes.